OpenStudy (.whitedragon.):

is the equation 3(2x-4)=-18 equivalent to 6 x-12=-18?

OpenStudy (anonymous):

Well you can figure that out

OpenStudy (anonymous):

First we need to simplify

OpenStudy (anonymous):

So for the first equation it would become 6x - 12 = -18

OpenStudy (anonymous):

aka 6x = -6 because you add the 12 to both sides to get rid of it

OpenStudy (anonymous):

aka x = -6/6

OpenStudy (.whitedragon.):

@ganeshie8

OpenStudy (anonymous):

which is equal to what?

OpenStudy (.whitedragon.):

-1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now we KNOW they are both the same without even finishing them because 3(2x -4) = 6x - 12

OpenStudy (anonymous):

Do you understand?

OpenStudy (.whitedragon.):

thanks hey is that ur real pic